### v2

# A Transformer Model for Predicting Chemical Reaction Products from Generic Templates

Abstract
The accurate prediction of chemical reaction outcomes is a major challenge in computational chemistry. Current models rely heavily on either highly specific reaction templates or template-free methods, both of which present limitations. To address these limitations, this work proposes the Broad Reaction Set (BRS), a dataset featuring 20 generic reaction templates that allow for the efficient exploration of the chemical space. Additionally, ProPreT5 is introduced, a T5 model tailored to chemistry that achieves a balance between rigid templates and template-free methods. ProPreT5 demonstrates its capability to generate accurate, valid, and realistic reaction products, making it a promising solution that goes beyond the current state-of-the-art on the complex reaction product prediction task.
I Introduction
The accurate prediction of chemical reaction outcomes is an important task in chemistry as it allows the construction of organic synthesis routes. Given a set of reactive molecules i.e. reactants, the goal is to determine the outcome, the product. This task is especially challenging, requiring a thorough understanding of chemical substances, compound classes, reactions, underlying reactivity patterns, and reaction conditions. Such an understanding is fundamental for various applications, including: Drug discovery [1], material science [2], and green chemistry [3].
Organic chemistry synthesis route planning has usually been an expert driven and rule based approach. However, the advancement of machine learning has transformed the field of cheminformatics. Among the most promising techniques in this field are Transformer models [4], originally developed for sequence-to-sequence tasks in Natural Language Processing (NLP) [5]. The Transformer architecture has revolutionized NLP by enabling models to understand context through self-attention mechanisms. This capability helps the models to learn to dynamically assign weights to different parts of the input data to produce more accurate and contextually relevant outputs. The Transformer architecture has also been successfully applied to the field of chemistry in various tasks, such as single-step chemical reaction prediction [6, 7, 8], retrosynthesis [8, 9, 10], molecule generation [11, 12, 13], and molecular property prediction[14, 15, 16]. However, these models are highly dependent on the quality and diversity of the data used for training.
Molecular data can be represented in the form of strings which are derived from a graph traversal of their molecular structure. This compact representation, known as the Simplified Molecular Input Line Entry System (SMILES) [17] encodes molecules where each character represents an atom or bond, providing a linear representation useful for computational purposes and database storage. When the graph traversal follows the specific order established by the notation, the resulting SMILES is canonical and a molecule has only one canonical SMILES but can have multiple non-canonical SMILES representations.
The SMILES notation can also be used to represent chemical reactions, encoding reactants (starting materials), reagents (substances that assist the reaction but do not transform), and products (final molecules produced). The SMARTS [18] notation, on the other hand, extends SMILES by representing patterns of atoms and bonds, or substructures within a molecule. It allows the identification of functional groups in reactants, creating reaction templates that capture general reactivity patterns and improve the prediction and recognition of chemical behavior.
The task of predicting reaction products studied in this paper is typically addressed using two different approaches: template-based methods, which rely on predefined rules (such as reactions represented by reaction SMILES or SMARTS), and template-free methods, which learn reactivity patterns directly from large datasets without relying on predefined rules or patterns.
Publicly available reaction datasets are mostly obtained through data extraction from patents published by The United States Patent and Trademark Office (USPTO) [19]. These reactions are represented using SMILES representation along with their reactants and reagents. Previous work has created a sub-dataset called USPTO MIT [20] by filtering it to 470,000 examples and splitting it into train/test/validation sets. USPTO MIT datasets are widely used in the literature.
Since USPTO datasets are derived from patents, they provide highly specific reactions tailored exclusively to particular reactant-product combinations. Furthermore, the chemical space explored using these datasets is limited to the information contained within the patented reactions. Consequently, we assume that relying solely on patents introduces bias during training, leaving a substantial portion of the chemical space unaccounted for. While previous work [8] has demonstrated great accuracy in predicting reaction outcomes using this dataset, the lack of molecule discovery and real-world applications stemming from such models highlights a key limitation: models trained solely on the patented chemical space suffer from generalization issues and are unsuitable for practical applications [21]. Additionally, the limitations of the existing datasets become particularly evident when reactions are considered as actions in reinforcement learning. The specificity of the reactions limits the exploration of all possible combinatorial space with these algorithms.
Such limitations and observations have led us to propose a new dataset called Broad Reaction Set (BRS) that bridges the gap between regular expressions and overly specific reaction SMILES. These reactions are expressed using the SMARTS syntax and represent broader transformation patterns rather than single, specific reactions. This approach enables us to explore a more comprehensive portion of the chemical space, which is a fundamental step for the tasks of organic synthesis and molecule discovery.
Reaction template datasets are often argued to be difficult to maintain due to the endless and unmanageable number of reactant-product combinations [7], leading many to favor template-free methods as the future of the field. While maintaining templates for highly specific reactions, such as those in USPTO, is undeniably challenging, we argue that employing generic reactions, like those in the proposed BRS, combines the benefits of reaction templates while addressing the maintenance issues.
To investigate the properties of BRS, we release a tailored T5 [22] model, ProPreT5, designed for the reaction prediction problem. ProPreT5 was separately trained in both template-based and template-free settings using the BRS and USPTO MIT datasets. This approach led to several key insights. When using reaction templates to predict products with the USPTO MIT dataset, the task proved to be trivial, as the product patterns in the reaction templates essentially enumerate the product molecule, significantly simplifying the prediction. The goal of this study is not to improve the USPTO MIT baseline but to highlight its limitations and propose a more realistic alternative. To this end, we opted for an economical training setup, using the USPTO MIT dataset as a sanity check for ProPreT5. Ultimately, we demonstrate that ProPreT5 generates realistic, valid, and accurate reaction products while allowing for the exploration of the entire chemical space in a template-based setting, addressing key challenges and going beyond the current state-of-the-art approaches in reaction prediction. An analysis aimed at improving the interpretability of the template-based ProPreT5 model is also proposed, leading to the conclusion that generic reaction templates provide the model with crucial information for product generation.
The contribution presented in this article is threefold:
  • We address the limitations of publicly available reaction datasets by introducing a novel reaction set BRS allowing for a broader exploration of the chemical space.
  • We release ProPreT5, an improved and highly flexible T5-based model capable of generating realistic, valid, and accurate products.
  • We perform an interpretability analysis of the model to identify key contextual information that affects the accuracy of product prediction.
TABLE I: Generic Reactions Patterns
The greater-than signs (>>much-greater-than$>$ >) separates reactants from products, while a dot (.) distinguishes individual molecules. #$\#n$ represents any atom with the atomic number $n$, and specific letters indicate particular chemical environments; $:n$ is used to map and track specific subgraphs within the reaction. For more information on the SMARTS notation refer to [18].
SMARTS Notation
# Constructive Reactions # Destructive Reactions
1 [#6,#7,#8;h:1].[O,N,F,C:2]>>[#6,#7,#8:1][O,N,F,C:2] 11 [#6,#7,#8:1][O,N,F,C:2]>>[#6,#7,#8;h:1]
2 [O,N,C;h:1][O,N,C;h:2]>>[O,N,C:1]=[O,N,C:2] 12 [O,N,C:1]=[O,N,C:2]>>[O,N,C;h:1][O,N,C;h:2]
3 [N,C;h2:1][N,C;h2:2]>>[N,C:1]#[N,C:2] 13 [N,C:1]#[N,C:2]>>[N,C;h2:1][N,C;h2:2]
4 [C;h:1]=[N,C;h:2]>>[C:1]#[N,C:2] 14 [C:1]#[N,C:2]>>[C;h:1]=[N,C;h:2]
5 [#6,#7,#8;h:1$\sim$[*:2$\sim$[#6,#7,#8;h:3]>> [#6,#7,#8:1]1[*:2$\sim$[#6,#7,#8:3]1 15 [#6,#7,#8:1]1[*:2$\sim$[#6,#7,#8:3]1>> [#6,#7,#8;h:1$\sim$[*:2$\sim$[#6,#7,#8;h:3]
6 [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[#6,#7,#8;h:3]>> [#6,#7,#8:1]1[*:2$\sim$[*:4$\sim$[#6,#7,#8:3]1 16 [#6,#7,#8:1]1[*:2$\sim$[*:4$\sim$[#6,#7,#8:3]1>> [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[#6,#7,#8;h:3]
7 [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5$\sim$[#6,#7,#8;h:3]>> [O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[#6,#7,#8:3]1 17 [O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[#6,#7,#8:3]1>> [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5$\sim$[#6,#7,#8;h:3]
8 [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$ [#6,#7,#8;h:3]>>[O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6] $\sim$ [#6,#7,#8:3]1 18 [O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$[#6,#7,#8:3]1 >>[#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$ [#6,#7,#8;h:3]
9 [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$[*:7$\sim$ [#6,#7,#8;h:3]>>[O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6] $\sim$[*:7$\sim$[#6,#7,#8:3]1 19 [O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$[*:7$\sim$ [#6,#7,#8:3]1>>[#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5] $\sim$[*:6$\sim$[*:7$\sim$[#6,#7,#8;h:3]
10 [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$[*:7$\sim$ [*:8$\sim$[#6,#7,#8;h:3]>>[O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5] $\sim$[*:6$\sim$[*:7$\sim$[*:8$\sim$[#6,#7,#8:3]1 20 [O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$[*:7$\sim$[*:8] $\sim$[#6,#7,#8:3]1>>[#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5] $\sim$[*:6$\sim$[*:7$\sim$[*:8$\sim$[#6,#7,#8;h:3]
II Related Work
Most of the recent literature on template-based approaches in reaction prediction has focused on graph-based models. [23, 24]. However, to the best of our knowledge, the potential of Transformer models for template-based sequence prediction has yet to be fully explored. This can be partly attributed to the specificity of reaction templates in the benchmark datasets (USPTO MIT).
Template-free, sequence-based models have achieved impressive results on the USPTO MIT benchmark, holding the current state-of-the-art in the single-step product prediction task [8]. However, these models face significant challenges in terms of applicability. The USPTO MIT dataset, limited to reactions found in patents, struggle to capture the diversity and complexity of real-world organic synthesis. As a result, while template-free models perform well in generating reaction outcomes, they struggle to generalize to the discovery of novel molecules or the prediction of reaction outcomes that fall outside the dataset [21].
Notable advancements in this domain include Molecular Transformer [6], which used a smaller version of the base transformers architecture [4]. The model was trained on both forward and backward tasks. This was the first application of transformers to the single-step reaction prediction task. Augmented Molformer [7] used data augmentation by extending the inputs with non-canonical SMILES and demonstrated that doing so increased the model’s generalization. Chemformer [8], on the other hand, took things even further by proposing a much larger BART [25] model with special pretraining. The pretraining included masking and data augmentation, making the model more robust and increasing prediction accuracy. While these template-free transformer models have demonstrated remarkable success in reaction prediction, the exploration of template-based sequence-to-sequence models for this task remains an underexplored area, leaving room for further development in this domain.
The lack of real-world applications using previous models highlights the need for a dataset with generic reactions, offering a more versatile training ground for reaction prediction models. This would allow models to explore the entire chemical space, ultimately advancing the field beyond the limitations of current benchmarks.
III Datasets
Two datasets are employed in this work: USPTO MIT [20] and the dataset constructed by using the proposed BRS which is one of the key contributions of this work. The USPTO MIT dataset contains reactions, along with the corresponding reactants, reagents, and products extracted from patents, represented in SMILES notation. The USPTO MIT dataset includes highly specific reaction templates, valid only for the exact combination of reactants, reagents, and products. It is divided into two subsets: USPTO MIT Mixed, which does not distinguish between reactants and reagents and is considered a slightly more challenging problem, and USPTO MIT Separated, which separates reagents from reactants.
In this study, we worked with the USPTO MIT Mixed dataset and performed a verification process using RDKit [26] to ensure consistency between reactants, reagents, products, and their corresponding reaction templates. With our specific configuration, we identified a small number of discrepancies where the combination of reactants, reagents, and products did not align with the reaction template. These mismatches indicated that the reactions could not be performed as defined, leading to the removal of 142 examples from the dataset. This adjustment had a negligible impact on the overall dataset size: the training set was reduced from 409,035 to 408,916 examples, the validation set from 30,000 to 29,990, and the test set from 40,000 to 39,987. The proposed train/validation/test split was preserved.
III-A Proposed Reaction Set: Broad Reaction Set
For the proposed dataset, 20 new generic reactions are introduced, as shown in Table I. Ten of these reactions are constructive, and the other ten are destructive, represented in SMARTS notation. These reactions are inspired by those used in an evolutionary algorithm named EvoMol, which demonstrated efficiency in exploring chemical space [27]. These reactions serve as foundational building blocks for constructing prediction datasets similar to USPTO. Since these reactions can be applied to a wide range of molecules, they offer significant flexibility for use with publicly available or commercial datasets.
Destructive reactions, the reverse of constructive reactions, enable a return to an earlier stage in the chemical space, offering the possibility to explore alternative pathways. Constructive and destructive reactions are symmetrical, and since this work focuses on single-step prediction, only constructive reactions are used in this study.
Here’s what each reaction from Table I does:
  • Reactions 1 & 11: The constructive reaction (#1) involves a molecule with an atom such as C, N, or O, bonded to hydrogen, reacting with a functional group containing O, N, F, or C. The functional group is added to the reactant. In contrast, the destructive reaction (#11) removes a functional group containing O, N, F, or C and replaces it with a hydrogen atom.
  • Reactions 2 & 12: The constructive reaction (#2) takes a molecule with a single bond between O, N, or C atoms which are bonded to H and forms a double bond instead. The destructive reaction (#12) breaks a double bond between those same heavy atoms, replacing it with a single bond.
  • Reactions 3 & 13: The constructive reaction (#3) takes a molecule with a single bond between N and C atoms, each bonded to two hydrogen atoms, and replaces this bond with a triple bond between these atoms, breaking the bond of one hydrogen atom for each. The destructive reaction (#13) breaks the triple bond and adds a hydrogen atom to each heavy atom that forms the bond.
  • Reactions 4 & 14: The constructive reaction (#4) takes a molecule with a C atom bonded to a H. This C atom contains a double bond with a N or C atom, which in turn is also bonded to at least one H. The reaction transforms the double bond into a triple bond. The destructive reaction (#14) takes a molecule with a triple bond between a C atom and a N or C atom and breaks this triple bond into a double bond.
The remaining reactions focus on the creation or destruction of cycles of various sizes and will be explained together:
  • Reactions 5 – 10  &  15 – 20: The constructive reactions (#5 – #10) involve a molecule with a linear structure where an atom C, N or O bonded to H is connected to one (for reaction 5) up to six (for reaction 10) wildcard atoms (any atom), forming a cyclic structure by bonding the atom at position :1 with the atom at position :3 to close the ring. The destructive reactions (#15 – #20) involve a molecule with a cycle of size 3 (for reaction 15) up to 8 (for reaction 20) and they break the cycle.
With their highly generic patterns, these reactions can be applied to a wide range of molecules, as well as different substructures within the same molecule. This makes the reaction set extremely flexible and well-suited for exploring chemical space. It is argued that these 20 reactions provide a solid foundation for starting with the simplest molecules and exploring a significant portion of the chemical space.
For simplicity, the transformations defined in these reactions are currently limited to atoms C, N, O, and F. However, this does not imply that the reactions cannot be applied to molecules containing other atoms. It simply means that only the reactivity of these four atoms is considered, and transformations will occur exclusively between them within the molecule. Additionally, reactions constructing cycles up to size 8 were defined. This limitation is partly due to challenges in defining reactions that can generate cycles of arbitrary size using SMARTS notation, but also because cycles larger than size 8 are rare. Moreover, chemists have proposed metrics that try to assess the synthesizability of a molecular graph such as SAScore [28] where large cycles are penalized.
Despite these limitations, the defined reactions allow us to explore a significant portion of the chemical space. With minor modifications, these reactions can be extended to include additional atoms and cover a larger part of the chemical space, exceeding the limits set by the current definition. For the time being, the limitations we have set still enable us to explore the chemical space.
III-B Construction of The Proposed Dataset
To construct the dataset used in this study, reactants were randomly sampled from several publicly available datasets: EVO10 [29], which enumerates every possible molecule with a maximum of 10 atoms of either C, N, O, F, or S; ZINC20 [30]; and ChEMBL34 [31], both of which were filtered to the same atoms as EVO10. For simplicity, the proposed dataset will be referred to as the BRS dataset, even though BRS is the name given to the reaction set employed for the construction of the dataset.
To create the dataset, molecules were randomly selected from the three datasets mentioned above while making sure that ZINC20, a dataset containing over 1B molecules, wasn’t overly represented. At each iteration, either a new molecule was selected from the datasets or a product from an earlier iteration was used as a reactant, allowing for the construction of multi-step synthesis routes for future studies. To avoid the generation of unrealistic products by the generic BRS, several filtering steps were applied. First, only canonical SMILES were kept. Additionally, filters described in [29] were implemented: one removed molecules with ECFP4 subgraphs absent from real-molecule datasets such as ChEMBL and ZINC, while the other excluded products containing Generic Cyclic Features (GCF)—scaffolds of molecular cycles not observed in the same datasets. Products that passed these filters were deemed realistic. To enhance diversity, multiple products were included for the same reactant-reaction combinations, allowing the model to learn that a single input could yield multiple valid outputs. We also ensured that each reaction appeared an equal number of times in the train, test, and validation datasets. The resulting train dataset contains 311,621 examples, while the validation and test datasets both contain 10,000 examples. Much larger datasets can be constructed by using the reactions from the BRS and by incorporating molecules from different datasets.
Figure 1: ProPreT5 Architecture.
IV Method
IV-A Model and Implementation Details
The general architecture of ProPreT5 is illustrated in Figure 1. A T5 model was chosen for its ease of use, relatively lightweight architecture, and ability to handle large datasets even with modest training resources. The model is very similar to the original T5 [22], encompassing an encoder-decoder structure. The encoder and decoder blocks each consisting of 6 layers with a hidden size of 512. Each block includes self-attention and feed-forward layers, with 8 attention heads. Feed-forward layers have an intermediate size of 2048 and use ReLU activation. In addition to these layers, the decoder includes cross-attention layers and masked self-attention. Relative positional embeddings are also used to capture token relationships.
Both the encoder and decoder share a vocabulary embedding layer with a vocabulary size of 243 tokens, as both the input and output use the same notation—SMILES for reactants and product and either SMILES or SMARTS for reactions. The relatively small vocabulary size results from character-level tokenization, which enhances ProPreT5’s flexibility. This approach allows most datasets to be used without retraining the model, unless a new character is added—unlikely given that the vocabulary covers a substantial portion of known chemistry. Character-level tokenization was selected to address the challenge of ensuring pre-trained models have the correct tokens, as adding and finetuning new tokens is costly. Additionally, a larger Byte-Pair Encoding (BPE) tokenizer [32] with around 10,000 tokens was trained and tested but this led to generalization issues. This compact and efficient model enables the generation of reaction outcomes.
As demonstrated in Figure 1, the reactants and the reaction are distinguished with a separation token used to separate both the reactants from one another and the reaction from the reactants. Additionally, type embeddings are applied to each input, following the approach in [11]. These embeddings distinguish between different input and output types, making it easier to incorporate new entry types, such as reagents, without confusing the model. A separate trainable embedding layer is used to map these type tokens to embedding vectors. ProPreT5 can also make predictions in a template-free setting, meaning the reaction templates would be excluded from the encoder input sequence.
TABLE II: Comparison of Reaction Product Prediction Accuracy
Comparisons were not provided for the template-based versions, as none of the models included the necessary tokens for reaction templates. Additionally, the Molecule Transformer code was unavailable, and we were unable to run the Augmented Transformer. (*) Results were obtained without fine-tuning. Results with citations were directly taken from the corresponding papers.
Dataset Reaction Template Molecule Transformer Augmented Transformer Chemformer ProPreT5
USPTO MIT Mixed Template-based - - - 99.8%
USPTO MIT Mixed Template-free 88.6% [6] 90.0% [7] 90.9% [8] 87.2%
BRS Template-based - - - 85.8%
BRS Template-free - - 5.6%* 8.8%*
IV-B Computational Resources and Training Setup
The training and evaluation of ProPreT5 were completed on a high-performance computing cluster equipped with NVIDIA V100 GPUs with 32 GB of memory, providing sufficient capacity to handle large datasets. ProPreT5 supports distributed training and was trained in parallel on 16 GPUs for 20 hours. Unlike some models in the literature, this setup was lightweight and time-efficient, yet even this relatively short 20-hour training provided satisfying results.
V Results and Discussion
ProPreT5 was trained in both template-free and template-based settings. The template-free training primarily served as a sanity check, allowing us to compare ProPreT5’s performance with existing models in the literature. This step was essential to confirm that the model was functioning as expected and could be reliably used with BRS. The goal was not to improve the USPTO MIT baseline but to ensure the model’s performance aligned reasonably well with prior work, despite using an economical training approach with minimal data augmentation. Specifically, we augmented the dataset by a factor of four using non-canonical SMILES as proposed in [33], where multiple non-canonical forms of reactants were paired with the same product. Additionally, costly pretraining methods, commonly employed in the existing literature, were not used in this case.
Table II presents the exact match accuracy of different models trained on various datasets. In the USPTO MIT Mixed dataset, each reactant-reaction combination corresponds to exactly one product, so if the model did not generate that specific product, the prediction was considered incorrect. In contrast, in the BRS dataset, a single reactant-reaction combination can lead to multiple possible products, similar to how multiple correct translations can exist for a machine translation task. Therefore, in the case of BRS, a generation was considered correct if the model predicted any one of the possible products.
V-A Template-Free Reaction Product Prediction
As seen in Table II, after only 20 epochs of training with a relatively economical setup, ProPreT5 achieved an exact match accuracy of 87.2% on the template-free USPTO MIT Mixed baseline. This result is particularly impressive considering the hundreds of epochs required by the other three models. This accuracy was sufficient to confirm that the model was functioning as expected and ready to tackle the new, ambitious task at hand.
Predicting reaction products for the BRS dataset is inherently more complex than for the USPTO dataset. Without the use of reaction templates, accurately predicting products becomes even more challenging. Since a single molecule can act as a reactant in many different reactions, the model faces significant limitations in predicting the correct transformation. As a result, the model can only learn general transformation patterns and apply them randomly, which makes it highly unlikely that the generated product matches the expected outcome. The results obtained without reaction templates for the BRS dataset highlight the complexity of this task in the absence of templates.
To provide a point of comparison, a test epoch was run on Chemformer using the publicly available code from [8], without any fine-tuning. Similarly, ProPreT5 was also tested without fine-tuning. We would have liked to offer the same comparison with the other two models, but the code for the Molecule Transformer was not publicly available, and we were unable to successfully run the code for the Augmented Transformer.
The low percentage of correct predictions by both models underscores the difficulty of this task and suggests that in real-world scenarios where a reactant can be involved in multiple different reactions, models trained in a template-free setting struggle to generalize and produce accurate predictions.
(a) Mean Aggregation of Importance Matrices
(b) Max Aggregation of Importance Matrices
Figure 2: Interpretability Analysis of Template-Based Generations of ProPreT5 on USPTO and BRS
(a) Reactant Tokens
(b) Reactant Template Tokens
(c) Product Template Tokens
Figure 3: Importance of Each Token in Generation for a Single Example from USPTO
V-B Template-Based Reaction Product Prediction
When it comes to template-based prediction, ProPreT5 achieves a near-perfect accuracy of 99.8% on the USPTO MIT Mixed baseline, suggesting that the task becomes trivial since the templates are highly specific and essentially enumerate the atoms in the product. In contrast, for the BRS dataset, the templates provide essential information necessary for generating the correct product. Without this information, making accurate predictions becomes an inherently illogical task. Given that the same reactants can participate in multiple reactions within BRS and yield different products, the model must effectively interpret the information encoded in the generic reaction templates to make accurate predictions. Despite this complexity, ProPreT5 achieves an impressive accuracy of 85.8%, highlighting its effectiveness even in more difficult prediction scenarios.
To confirm the intuition about template-based predictions on USPTO and BRS, we focused on understanding how the model utilizes the information in the input sequence. For the analysis, we used Inseq [34], a tool for attributing importance scores to input tokens, which helps understand how the model generates its outputs. Integrated gradients were used as the attribution method. This analysis was conducted to enhance the interpretability of ProPreT5.
In Figure 2, the template-based generations were analyzed. The input sequence was divided into three subcategories: Reactants, the reactant template, and the product template of the reaction (typically separated by >>much-greater-than$>$ > in SMARTS notation). As a result, three importance score matrices were obtained using Inseq for each example, aligning with each of the three input sections. These matrices were either max-aggregated or mean-aggregated to provide a general overview of the entire test dataset. The violin plots in Figure 2 display the distribution of mean and max importance scores for the template-based predictions on the BRS and USPTO datasets.
As can be deduced by looking at both Figure 2(a) and Figure 2(b), most of the information in the input sequence is not utilized for generation on USPTO. Only a few tokens carry the essential information. These results support the earlier hypothesis that when generating a product, the model primarily focuses on a handful of tokens from the reactant part and the product template part of the input. The model focuses mainly on specific tokens, while the rest of the input contributes little to the generation.
Figure 2 shows the entire dataset, while Figure 3 highlights an example where darker colors indicate higher contribution to the output, and lighter colors indicate less contribution. This example from USPTO demonstrates that the contribution of the reactant pattern is minimal, as the target product is almost explicitly encoded in the reaction SMILES template. The model doesn’t need to find an alignment between the template reactants and the actual reactants to produce the result. This explains the near-perfect accuracy of 99.8% and why the task is trivial.
In contrast, the BRS dataset presents a more implicit relationship between the product and the reaction template. The mean aggregated importance graphs in Figure 2(a) for both parts of the reaction template are nearly identical, highlighting the model’s need to rely equally on both parts to make a correct prediction.
V-C Assesment Metrics for Template-Based ProPreT5
The template-based ProPreT5 model proved itself to be a good compromise between overly specific templates and template-free approaches. Its performance was further evaluated across various metrics, as presented in Table III.
Accuracy, as previously defined, refers to the exact match accuracy. For the BRS dataset, where multiple products can result from a single reactant-reaction combination, a prediction is considered correct if the model generates any one of the possible outcomes. Accuracy is then calculated as the ratio of correct predictions to the total number of predictions.
Validity evaluates whether the predicted molecules can be successfully transformed into molecular graphs using RDKit [26]. If a molecule violates the fundamental construction rules of organic chemistry, it cannot be parsed by RDKit and is thus considered invalid. Validity is determined as the ratio of parsable molecules to the total number of predictions.
On the other hand, RDKit isn’t capable of checking the realism of a product, and just because a molecule is valid doesn’t necessarily mean it’s realistic. Realism assesses whether the predicted molecules resemble known real-world molecules. This is achieved using subgraphs filters [29], which eliminate molecules containing groups or cycles never encountered in datasets of billions of real molecules. Realism is calculated as the proportion of predictions that pass these filters out of the total number of predictions.
The generic reactions from BRS are capable of generating valid but unrealistic products. As explained earlier in Section III-B, we filtered out those unrealistic products to construct the dataset. Table III demonstrates that ProPreT5, trained on this dataset, has the ability to generate valid, accurate, and realistic products, highlighting its utility in exploring and expanding the chemical space while respecting the rules of organic chemistry.
TABLE III: Metrics for Template-Based ProPreT5
Accuracy Validity Realism
85.8% 99.3% 91.0%
VI Conclusion
In this study, we introduced the Broad Reaction Set (BRS), a novel generic reaction set designed to provide a more realistic and versatile benchmark. This important contribution addresses key limitations of the widely used USPTO dataset. Until now, template-based reaction prediction models relied on overly specific reaction templates, limiting their real-world applicability that we seek in cheminformatics.
The relevance of the BRS dataset was supported by the presented ProPreT5, a flexible reaction product prediction model capable of generating valid, realistic, and accurate products. Its flexibility comes not only from training on a dataset with generic reactions but also from its character-level tokenization, which enables easy adaptation to other datasets without costly retraining processes. ProPreT5 is a major contribution as it differs from existing models by predicting products in a way that can be applied to real-world chemical problems. We also demonstrated that obtaining satisfactory results in lighter training setups is possible.
Future work will move beyond single-step product predictions and explore multi-step organic synthesis routes while continuing to use a template-based approach with ProPreT5. Additionally, we will focus on improving the prediction accuracy to avoid propagating the error in a multi-step setup.
Acknowledgment
The Acknowledgment section will be left empty at this time in order to maintain the anonymity of the submission.
Code and Data Availability
The code and the data for this work will be made available in the future as they cannot be shared at this time to maintain the anonymity of the submission.